OpenStudy (anonymous):

A person’s approximate brain weight is directly proportional to his or her body weight, and a person weighing 150 lb has an approximate brain weight of 4lb. a. Find a mathematical model expressing the approximate brain weight of a person as a function of the person’s body weight. b. Find the approximate brain weight of a person whose body weight is 176 lb.

OpenStudy (texaschic101):

(a)sorry...not good with functions (b) 150/4 = 176/x (150 lbs to 4 lbs = 176 lbs to x lbs) cross multiply (150)(x) = (4)(176) 150x = 704 x = 704/150 x = 4.69 lbs.

OpenStudy (anonymous):

its ok :)